OpenStudy (nikeboi101):

What is____when z = −3?

OpenStudy (nikeboi101):

hold on

OpenStudy (nikeboi101):

this is my last 1

pooja195 (pooja195):

post it :)

pooja195 (pooja195):

Simply plug in the given z value 4(-3)-3/-3+6 4 * -3=?

OpenStudy (nikeboi101):

ok

pooja195 (pooja195):

4 * -3= ?

OpenStudy (nikeboi101):

-12

pooja195 (pooja195):

Good -12-3=?

OpenStudy (nikeboi101):

-9

pooja195 (pooja195):

When we subtract a number from a negative the number will increase (negatively) try again -12-3=?

OpenStudy (nikeboi101):

oh 15

OpenStudy (nikeboi101):

-15

pooja195 (pooja195):

Yes -15 :-) now for the bottom -3+6=?

OpenStudy (nikeboi101):

3

pooja195 (pooja195):

Good -15/3=?

OpenStudy (nikeboi101):

-3

pooja195 (pooja195):

-15 divided by 3= ?

OpenStudy (nikeboi101):

oh i said 3 srry -5

pooja195 (pooja195):

Correct :)